# SEA: Self-Evolution Agent with Step-wise Reward for Computer Use

### v1

@@ III-B Filtering data with step model
For the step filtering model, its task is to determine whether the action has been successfully executed based on the two screenshots before and after this step, the thought and action information. We manually annotated some ground truth data and continuously adjusted prompts to ensure that the model’s performance follows our filtering criteria. Then, we randomly extracted 10,000 step data from the task trajectory and used the QWen2.5-VL-72B [1] model and adjusted the prompt for inference. Each piece of data was subjected to multiple inferences and voting to ensure accuracy as much as possible, and finally used to train the step filtering model. Through this method, we distill the capabilities of the large model into the small model, making the step filtering model lighter while maintaining excellent performance consistent with the large model.
Figure 4: Illustration of Trajectory Reasoning by Step-wise Reinforcement Learning.
III-C Step-wise Reinforcement Learning
@@
In this part, we will introduce Trajectory Reasoning by Step-wise Reinforcement Learning. Specifically, we will use the UI-TARS-7B [38] as our cold-start model, and based on this model, we will use the proposed methods to produce the final agent.
III-C1 Preliminary of GRPO
@@
Although there are lots of improvements in the GRPO, in this paper, we still choose it, as a result of its balance between performance and tidiness. Group Relative Policy Optimization (GRPO) [31] is a reinforcement learning method designed to the efficient fine-tuning of large language models (LLMs) without relying on an explicit value function or critic. GRPO replaces the need for a trained value function in standard Proximal Policy Optimization (PPO) with the average rewards of multiple outputs for the same problem, making it particularly suitable for LLMs.
Given a batch of $G$ responses $\{o_{i}\}_{i=1}^{G}$ from a query $q$, each consisting of a sequence of tokens $o_{i}=(o_{i}(1),\dots,o_{i}(T))$, the GRPO objective is defined as:
$\mathcal{J}_{\text{GRPO}}(\theta)$ $=\mathbb{E}[\begin{subarray}{c}q\sim\mathcal{P}(Q),\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(O|q)\end{subarray}]$ (1)
$\quad\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\Bigg{\{}\min\Bigg{[}\frac{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta_{\text{old}}}(o_{i,t}|q,o_{i,<t})}\hat{A}_{i,t},$
@@
$\qquad\text{clip}\left(\frac{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta_{\text{old}}}(o_{i,t}|q,o_{i,<t})},1-\varepsilon,1+\varepsilon\right)\hat{A}_{i,t}\Bigg{]}$
@@
$\qquad-\beta\mathbb{D}_{\text{KL}}\left[\pi_{\theta}\|\pi_{\text{ref}}\right]\Bigg{\}}$
Where $\mathbb{D}_{\text{KL}}\left[\pi_{\theta}\|\pi_{\text{ref}}\right]$ is shown in the following equation 2. $\hat{A}_{i,t}$ is the group-normalized advantage for token t in response $o_{i}$, computed as equation 3.
$\mathbb{D}_{\text{KL}}\left[\pi_{\theta}\|\pi_{\text{ref}}\right]=\frac{\pi_{\text{ref}}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}-\log\frac{\pi_{\text{ref}}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}-1$ (2)
$\hat{A}_{i,t}=\frac{r_{i}-\mu}{\sigma},\quad\text{with }r_{i}\text{ the total reward of }o_{i}$ (3)
@@
Where $\mu$ and $\sigma$ are the mean and standard deviation of rewards in the group.
III-C2 Step-wise RL training
@@ III-C3 Rewards
III-C4 Training Objective
We divide each task trajectory $\tau=(s_{0},a_{0},\ldots,s_{n},a_{n})$ into multiple steps. At each step, the agent observes the screen state $s_{i}$ and generates an action $a_{i}$ to fulfill the task instruction $x_{i}\in D$. Our training objective is to maximize the expected reward for step completion, consistency between thoughts and actions, and action format:
$\max_{\theta}\mathbb{E}_{x_{i}\sim D,\tau_{i}\sim\pi_{\theta}}[r_{i}(x_{i},\tau_{i})+r_{c}(x_{i},\tau_{i})+r_{f}(x_{i},\tau_{i})]$ (4)
@@ III-C4 Training Objective
III-D Inference Stage
During the inference stage, we build upon the GTA1 [41] and replace the original inference model with our proposed model SEA. As a result, both the sampling and selection of candidate trajectories during inference are handled entirely by SEA, without relying on any external scoring models. This design significantly reduces computational overhead and effectively improves decision-making efficiency.
TABLE I: Comparison of Various Planners and Grounding Methods on ScreenSpot-Pro.
@@ III-D Inference Stage
QwenVL-7B [42] 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.7 0.0 0.4 0.0 0.0 0.0 0.0 0.0 0.0 0.1 0.0 0.1
GPT-4o [43] 1.3 0.0 0.7 1.0 0.0 0.6 2.0 0.0 1.5 2.1 0.0 1.2 1.1 0.0 0.9 0.0 0.0 0.0 1.3 0.0 0.8
SeeClick [23] 0.6 0.0 0.3 1.0 0.0 0.6 2.5 0.0 1.9 3.5 0.0 2.0 1.1 0.0 0.9 2.5 0.0 1.5 1.8 0.0 1.1
Qwen2-VL-7B [44] 2.6 0.0 1.3 1.5 0.0 0.9 0.5 0.0 0.4 6.3 0.0 3.5 3.4 1.9 3.0 0.9 0.0 0.5 2.5 0.2 1.6
OS-Atlas-4B [45] 7.1 0.0 3.7 3.0 1.4 2.3 2.0 0.0 1.5 9.0 5.5 7.5 5.1 3.8 4.8 5.6 0.0 3.1 5.0 1.7 3.7
ShowUI-2B [46] 16.9 1.4 9.4 9.1 0.0 5.3 2.5 0.0 1.9 13.2 7.3 10.6 15.3 7.5 13.5 10.3 2.2 6.6 10.8 2.6 7.7
CogAgent-18B [47] 14.9 0.7 8.0 9.6 0.0 5.6 7.1 3.1 6.1 22.2 1.8 13.4 13.0 0.0 10.0 5.6 0.0 3.1 12.0 0.8 7.7
Aria-UI [48] 16.2 0.0 8.4 23.7 2.1 14.7 7.6 1.6 6.1 27.1 6.4 18.1 20.3 1.9 16.1 4.7 0.0 2.6 17.1 2.0 11.3
UGround-7B [49] 26.6 2.1 14.7 27.3 2.8 17.0 14.2 1.6 11.1 31.9 2.7 19.3 31.6 11.3 27.0 17.8 0.0 9.7 25.0 2.8 16.5
Claude Computer Use 22.0 3.9 12.6 25.9 3.4 16.8 14.5 3.7 11.9 33.9 15.8 25.8 30.1 16.3 26.9 11.0 4.5 8.1 23.4 7.1 17.1
OS-Atlas-7B [45] 33.1 1.4 17.7 28.8 2.8 17.9 12.2 4.7 10.3 37.5 7.3 24.4 33.9 5.7 27.4 27.1 4.5 16.8 28.1 4.0 18.9
UGround-V1-7B [49] 35.5 27.8 13.5 38.8 48.8 26.1 31.1
UI-TARS-2B [38] 47.4 4.1 26.4 42.9 6.3 27.6 17.8 4.7 14.6 56.9 17.3 39.8 50.3 17.0 42.6 21.5 5.6 14.3 39.6 8.4 27.7
UI-TARS-7B [38] 58.4 12.4 36.1 50.0 9.1 32.8 20.8 9.4 18.0 63.9 31.8 50.0 63.3 20.8 53.5 30.8 16.9 24.5 47.8 16.2 35.7
UI-TARS-72B [38] 63.0 17.3 40.8 57.1 15.4 39.6 18.8 12.5 17.2 64.6 20.9 45.7 63.3 26.4 54.8 42.1 15.7 30.1 50.9 17.5 38.1
@@
III-E Grounding-Based Generalization Enhancement.
@@
III-E1 Train Grounding Model
@@
Aguvis [50] and OS-Atlas[45] have organized and filtered many previous open-source datasets with high quality, so we have conducted our data filtering based on them. We select 30000 pieces of data from each subset of these two datasets (if fewer than 30000 pieces, we take all), and these 30000 pieces of data should come from different screenshots, rather than other elements of the same screenshot. We used these different datasets and the same configuration for SFT training, and evaluated them using screenspot-v2. Finally, we selected five higher-quality subsets from two large datasets as the training set for the grounding model. Through these experiments, we also found that the quality of the dataset is far more critical than the quantity.
In addition, detailed instructions are helpful for the grounding task. To reconstruct the data samples, we retain the first two steps from [51] For sample construction, we retain the first two steps from [51] while optimizing the third step. The original method discards samples with imprecise predicted coordinates during sample construction. However, we observe that these discarded samples often correspond to more complex tasks, involving richer operational logic and reasoning paths. Therefore, we preserve the original task instructions (the original commands) of these samples and reincorporate them into the training set. To further enhance data diversity and model robustness, we apply data augmentation techniques such as image scaling and color inversion to the filtered dataset.
@@
$\text{Reward}=\min\Bigg{(}1,$ $\text{IoU}(B_{p},B_{gt})\cdot$ (5)
@@
III-E2 Model Merge
After getting the grounding model, we need to find a simple and effective way to combine the planning and grounding abilities into one model. Meanwhile, we do not want to use extra computation and data to distill these two models into one. As a result of that, we leverage the DARE [52] to merge these two models. Based on the idea of task vectors, this method sets most incremental parameters (i.e., the differences between parameters of the grounding model and planning model) to zero without affecting the performance of original models. DARE randomly discards incremental parameters at a rate of $p$ and rescales the remaining parameters by 1/(1p)1/(1-p) to approximate the original embedding.
III-F Temporal Compressed Sensing Mechanism

### v2

@@ 3.1.3. Filtering data with step model
For the step filtering model, its task is to determine whether the action has been successfully executed based on the two screenshots before and after this step, the thought and action information. We manually annotated some ground truth data and continuously adjusted prompts to ensure that the model’s performance follows our filtering criteria. Then, we randomly extracted 10,000 step data from the task trajectory and used the Qwen2.5-VL-72B (Shuai et al., 2025) model and adjusted the prompt for inference. Each piece of data was subjected to multiple inferences and voting to ensure accuracy as much as possible, and finally used to train the step filtering model. Through this method, we distill the capabilities of the large model into the small model, making the step filtering model lighter while maintaining excellent performance consistent with the large model.
Figure 4. Illustration of Trajectory Reasoning by Step-wise Reinforcement Learning. This figure illustrates the process of trajectory reasoning via step-wise reinforcement learning (RL). Step-wise RL allows an agent to operate and obtain rewards at each step from Step 0 to Step n along a task trajectory. Each step comprises Input, Thought, Action, and Ground Truth components, with rewards calculated based on four facets: Overall Reward, Consistency Reward, Format Reward, and Step Reward.
3.2. Step-wise Reinforcement Learning
@@
In this part, we will introduce Trajectory Reasoning by Step-wise Reinforcement Learning. Specifically, we will use the UI-TARS-7B (Yujia et al., 2025) as our cold-start model, and based on this model, we will use the proposed methods to produce the final agent.
3.2.1. Preliminary of GRPO
Although there are lots of improvements in the GRPO, in this paper, we still choose it, as a result of its balance between performance and tidiness.
@@ 3.2.1. Preliminary of GRPO
Group Relative Policy Optimization (GRPO) (Zhihong et al., ) is a reinforcement learning method designed for efficient fine-tuning of large language models (LLMs) without relying on an explicit value function or critic. GRPO replaces the trained value function in standard Proximal Policy Optimization (PPO) with the average rewards of multiple outputs for the same query, making it particularly suitable for LLMs. Given a batch of $G$ responses $\{o_{i}\}_{i=1}^{G}$ from a query $q$, each consisting of a sequence of tokens $o_{i}=(o_{i}(1),\dots,o_{i}(T))$, the GRPO objective is defined as:
(1) $\mathcal{T}_{\text{GRPO}}(\theta)$ $=\mathbb{E}\left[q\sim\mathcal{P}(Q),\left\{o_{i}\right\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(O|q)\right]$
[…]
@@
$\quad\left.\text{clip}\left(\frac{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta_{\text{old}}}(o_{i,t}|q,o_{i,<t})},1-\varepsilon,1+\varepsilon\right)\hat{A}_{i,t}\right]$
@@
$\quad\left.-\beta\mathbb{D}_{\text{KL}}\left[\pi_{\theta}\parallel\pi_{\text{ref}}\right]\right\}$
where $\mathbb{D}_{\text{KL}}[\pi_{\theta}\parallel\pi_{\text{ref}}]$ denotes the Kullback-Leibler divergence between the current policy $\pi_{\theta}$ and the reference policy $\pi_{\text{ref}}$, defined in Equation 1:
(2) $\mathbb{D}_{\text{KL}}\left[\pi_{\theta}\parallel\pi_{\text{ref}}\right]=\frac{\pi_{\text{ref}}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}-\log\frac{\pi_{\text{ref}}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}-1$
$\hat{A}_{i,t}$ represents the group-normalized advantage for token $t$ in response $o_{i}$, calculated as:
(3) $\hat{A}_{i,t}=\frac{r_{i}-\mu}{\sigma}$
@@
Where the $r_{i}$ is the total reward of $o_{i}$, while $\mu$ and $\sigma$ are the mean and standard deviation of rewards in the group.
3.2.2. Step-wise RL training
@@ 3.2.3. Rewards
3.2.4. Training Objective
We divide each task trajectory $\tau=(s_{0},a_{0},\allowbreak\ldots,\allowbreak s_{n},a_{n})$ into multiple steps. At each step, the agent observes the screen state $s_{i}$ and generates an action $a_{i}$ to fulfill the task instruction $x_{i}\in D$. Our training objective is to maximize the expected reward for step completion, consistency between thoughts and actions, and action format:
(4) $\max_{\theta}\mathbb{E}_{x_{i}\sim D,\tau_{i}\sim\pi_{\theta}}[r_{i}(x_{i},\tau_{i})+r_{c}(x_{i},\tau_{i})+r_{f}(x_{i},\tau_{i})]$
@@ 3.2.4. Training Objective
Table 1. Comparison of Various Planners and Grounding Methods on ScreenSpot-Pro. It presents a comparison of various planners and grounding methods in terms of their performance on the ScreenSpot-Pro dataset. It reports results across different task categories (Development, Creative, CAD, Scientific, Office, and OS) for both Text and Icon modalities, along with the average (Avg) scores for each agent model.
@@
QwenVL-7B (Bai et al., ) 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.7 0.0 0.4 0.0 0.0 0.0 0.0 0.0 0.0 0.1 0.0 0.1
GPT-4o (Hurst et al., 2024) 1.3 0.0 0.7 1.0 0.0 0.6 2.0 0.0 1.5 2.1 0.0 1.2 1.1 0.0 0.9 0.0 0.0 0.0 1.3 0.0 0.8
SeeClick (Kanzhi et al., 2024) 0.6 0.0 0.3 1.0 0.0 0.6 2.5 0.0 1.9 3.5 0.0 2.0 1.1 0.0 0.9 2.5 0.0 1.5 1.8 0.0 1.1
Qwen2-VL-7B (Wang et al., 2024a) 2.6 0.0 1.3 1.5 0.0 0.9 0.5 0.0 0.4 6.3 0.0 3.5 3.4 1.9 3.0 0.9 0.0 0.5 2.5 0.2 1.6
OS-Atlas-4B (Zhiyong et al., 2024b) 7.1 0.0 3.7 3.0 1.4 2.3 2.0 0.0 1.5 9.0 5.5 7.5 5.1 3.8 4.8 5.6 0.0 3.1 5.0 1.7 3.7
ShowUI-2B (Qinghong et al., 2025) 16.9 1.4 9.4 9.1 0.0 5.3 2.5 0.0 1.9 13.2 7.3 10.6 15.3 7.5 13.5 10.3 2.2 6.6 10.8 2.6 7.7
CogAgent-18B (Hong et al., 2024) 14.9 0.7 8.0 9.6 0.0 5.6 7.1 3.1 6.1 22.2 1.8 13.4 13.0 0.0 10.0 5.6 0.0 3.1 12.0 0.8 7.7
Aria-UI (Yuhao et al., 2024) 16.2 0.0 8.4 23.7 2.1 14.7 7.6 1.6 6.1 27.1 6.4 18.1 20.3 1.9 16.1 4.7 0.0 2.6 17.1 2.0 11.3
UGround-7B (Gou et al., 2024) 26.6 2.1 14.7 27.3 2.8 17.0 14.2 1.6 11.1 31.9 2.7 19.3 31.6 11.3 27.0 17.8 0.0 9.7 25.0 2.8 16.5
Claude Computer Use (Hu et al., ) 22.0 3.9 12.6 25.9 3.4 16.8 14.5 3.7 11.9 33.9 15.8 25.8 30.1 16.3 26.9 11.0 4.5 8.1 23.4 7.1 17.1
OS-Atlas-7B (Zhiyong et al., 2024b) 33.1 1.4 17.7 28.8 2.8 17.9 12.2 4.7 10.3 37.5 7.3 24.4 33.9 5.7 27.4 27.1 4.5 16.8 28.1 4.0 18.9
UGround-V1-7B (Gou et al., 2024) 35.5 27.8 13.5 38.8 48.8 26.1 31.1
UI-TARS-2B (Yujia et al., 2025) 47.4 4.1 26.4 42.9 6.3 27.6 17.8 4.7 14.6 56.9 17.3 39.8 50.3 17.0 42.6 21.5 5.6 14.3 39.6 8.4 27.7
UI-TARS-7B (Yujia et al., 2025) 58.4 12.4 36.1 50.0 9.1 32.8 20.8 9.4 18.0 63.9 31.8 50.0 63.3 20.8 53.5 30.8 16.9 24.5 47.8 16.2 35.7
UI-TARS-72B (Yujia et al., 2025) 63.0 17.3 40.8 57.1 15.4 39.6 18.8 12.5 17.2 64.6 20.9 45.7 63.3 26.4 54.8 42.1 15.7 30.1 50.9 17.5 38.1
@@
Table 2. Grounding results on the ScreenSpot-V2. It presents the grounding results of various models on the ScreenSpot-V2 dataset, including their performance in Text and Icon modalities across three platforms (Desktop, Mobile, Web) as well as the average (AVG) scores.
MODEL Desktop Mobile Web AVG
Text Icon Text Icon Text Icon
UI-Venus-72B (Gu et al., 2025) 95.9 90.0 99.7 93.8 96.2 92.6 95.3
Holo1.5-72B (Company, 2025) 96.1 86.9 100.0 92.7 95.9 90.7 94.4
UI-Venus-7B (Gu et al., 2025) 96.9 90.7 99.0 90.0 96.2 88.7 94.1
Holo1.5-7B (Company, 2025) 96.7 88.9 99.2 91.1 95.4 86.1 93.3
GUI-ARP-7B (Ye et al., 2025) 97.2 85.7 96.5 89.0 94.3 84.8 91.8
Holo1.5-3B (Company, 2025) 95.0 89.7 99.2 88.0 91.8 84.8 91.7
Qwen2.5-VL-7B-Instruct (Shuai et al., 2025) 87.6 65.7 99.0 84.4 90.2 79.8 86.5
AGUVIS-7B (Yiheng et al., 2024) 93.3 77.9 95.5 81.5 91.0 77.8 87.3
OSAtlas-7B (Zhiyong et al., 2024b) 90.7 63.6 93.8 73.9 89.7 77.3 83.3
ZonUI-3B (Hsieh et al., ) 92.3 74.3 98.6 82.9 88.0 74.4 86.6
UGround-7B (Gou et al., 2025) 85.1 61.4 84.5 61.6 84.6 71.9 76.3
ShowUI-2B (Qinghong et al., 2025) 78.9 59.3 92.1 75.4 84.2 61.1 77.3
OSAtlas-4B (Zhiyong et al., 2024b) 64.4 46.4 82.8 64.0 78.6 60.1 68.5
CogAgent-18B (Hong et al., 2024) 75.8 20.7 69.3 27.0 74.4 31.5 52.8
MiniCPM-V-7B (Yuan et al., 2024) 69.1 18.6 59.7 18.0 67.9 26.6 45.9
SeeClick-7B (Kanzhi et al., 2024) 69.6 30.7 77.9 48.3 57.3 22.2 53.9
Qwen2.5-VL-3B-Instruct (Shuai et al., 2025) 54.1 30.0 62.1 46.4 31.2 48.3 46.9
GPT-4o (Hurst et al., 2024) 24.2 19.3 26.6 24.2 12.8 11.8 20.1
Ours 98.3 84.8 94.3 77.1 91.0 84.7 89.6
Table 3. Results with max steps of 50 on the online OS-World Benchmark. We have submitted our final model to the OSWorld website and compared with other SOTA methods, and our method achieves the best performance with a small number of parameters.
Category Model Name Succ Rate Chrome Gimp Libre Calc Libre Impress Libre Writer Multi Apps OS Thunder bird VLC VS Code
Closed Source Model claude-4-sonnet-20250514 $43.9$ $24.96$ $13.00$ $15.00$ $21.96$ $14.00$ $26.50$ $11.00$ $11.00$ $7.00$ $14.00$
claude-3-7-sonnet-20250219 $35.8$ $23.96$ $10.00$ $15.00$ $16.96$ $10.00$ $16.42$ $12.00$ $8.00$ $4.00$ $13.00$
o3 $17.2$ $10.00$ $10.00$ $4.00$ $2.00$ $5.00$ $11.00$ $9.00$ $3.00$ $2.99$ $5.00$
computer-use-preview $31.3$ $16.96$ $9.00$ $7.00$ $13.96$ $6.00$ $14.70$ $17.00$ $10.00$ $2.00$ $16.00$
Large Params opencua-32b (Wang et al., 2025) $34.1$ $18.96$ $18.00$ $7.00$ $15.99$ $10.00$ $12.35$ $12.00$ $8.00$ $5.00$ $13.00$
uitars-72b-dpo (Yujia et al., 2025) $25.8$ $14.96$ $16.00$ $6.00$ $11.96$ $10.00$ $6.24$ $8.00$ $5.00$ $4.00$ $11.00$
Small Params opencua-7b (Wang et al., 2025) $28.2$ $17.96$ $10.00$ $5.00$ $16.96$ $8.00$ $8.69$ $10.00$ $6.00$ $5.44$ $11.00$
uitars-1.5-7b (Seed, 2025) $27.3$ $12.96$ $13.00$ $2.00$ $16.99$ $9.00$ $9.09$ $6.00$ $7.00$ $3.00$ $11.00$
opencua-qwen2-7b (Wang et al., 2025) $20.6$ $14.96$ $6.00$ $3.00$ $12.00$ $10.00$ $4.00$ $2.00$ $7.00$ $5.00$ $10.00$
opencua-a3b (Wang et al., 2025) $19.9$ $9.92$ $14.00$ $1.00$ $10.96$ $8.00$ $6.07$ $9.00$ $1.00$ $2.00$ $10.00$
Ours $29.8$ $18.96$ $15.00$ $3.00$ $17.99$ $13.00$ $5.58$ $8.00$ $6.00$ $5.83$ $16.00$
3.3. Grounding-Based Generalization Enhancement.
@@
3.3.1. Train Grounding Model
@@
Aguvis (Yiheng et al., 2024) and OS-Atlas(Zhiyong et al., 2024b) have organized and filtered many previous open-source datasets with high quality, so we have conducted our data filtering based on them. We select 30000 pieces of data from each subset of these two datasets (if fewer than 30000 pieces, we take all), and these 30000 pieces of data should come from different screenshots, rather than other elements of the same screenshot. We used these different datasets and the same configuration for SFT training, and evaluated them using screenspot-v2. Finally, we selected five higher-quality subsets from two large datasets as the training set for the grounding model. Through these experiments, we also found that the quality of the dataset is far more critical than the quantity.
In addition, detailed instructions are helpful for the grounding task. To reconstruct the data samples, we retain the first two steps from (Zongru et al., 2025) For sample construction, we retain the first two steps from (Zongru et al., 2025) while optimizing the third step. The original method discards samples with imprecise predicted coordinates during sample construction. However, we observe that these discarded samples often correspond to more complex tasks, involving richer operational logic and reasoning paths. Therefore, we preserve the original task instructions (the original commands) of these samples and reincorporate them into the training set. To further enhance data diversity and model robustness, we apply data augmentation techniques such as image scaling and color inversion to the filtered dataset.
@@
(5) $\text{Reward}=\min\Bigg(1,$ $\text{IoU}(B_{p},B_{gt})\cdot$
@@
3.3.2. Model Merge
After getting the grounding model, we need to find a simple and effective way to combine the planning and grounding abilities into one model. Meanwhile, we do not want to use extra computation and data to distill these two models into one. As a result of that, we leverage the DARE (Yu et al., 2024) to merge these two models. Based on the idea of task vectors, this method sets most incremental parameters (i.e., the differences between parameters of the grounding model and planning model) to zero without affecting the performance of original models. DARE randomly discards incremental parameters at a rate of $p$ and rescales the remaining parameters by 1/(1p)1/(1-p) to approximate the original embedding.
3.3.3. Temporal Compressed Sensing Mechanism